# MIST: Towards Multi-dimensional Implicit BiaS Evaluation of LLMs for Theory of Mind

### v2

@@ B.1 Lexicon Selection
| (Low Morality) | Unprincipled, Unethical, Dishonest, Deceptive, Unaccountable, Corrupt, Unreliable, Duplicitous, Hypocritical, Negligent, Unscrupulous, Fraudulent, Deceitful, Manipulative, Unjust, Biased, Lawless, Self-serving, Exploitative, Opportunistic |
B.2 Data Construction Templates
Word Association Bias Test.
For WABT task, we use color coding to distinguish different categories of words for clarity in the data templates. Specifically, green tags indicate either the advantaged group identifiers () or positive attribute words (), while red tags indicate either the disadvantaged group identifiers () or negative attribute words ().
Affective Attribution Test.
In addition, blue tags are used to represent neutral object words () as well as group identity placeholders (), which may refer to either advantaged or disadvantaged social groups depending on the context for AAT task.
B.3 Data Statistics
For the WABT task we have 4,500 instances, and for the AAT task we have 1,500 instances. The detailed distribution of dataset across Race, Gender and Health domains is shown in Figure 6.
| Number | Percentage | |
| Total | 6000 | - |
| -WABT | 4500 | 75% |
| -AAT | 1500 | 25% |
| Domain | - | - |
| -Race | 2400 | 40% |
| -Gender | 1200 | 20% |
| -Health | 2400 | 40% |
B.4 Data Examples
In this section, we provide several concrete data examples from WABT and AAT, as shown in Figure 7. These examples cover different bias dimensions, including race, gender, and health. Each box presents representative input instances.
Appendix C Model Responses
C.1 Word Association Bias Test
The example C.1 is a valid paired example following the requirement. This example C.1 illustrates an instance of an invalid pairing. As observed, the model does not directly respond with a clear pairing strategy; instead, it adopts a seemingly random approach to avoid providing a definitive answer.
C.2 Affective Attribution Test
The example C.2 is a valid instance of AAT. The example C.2 is an invalid instance of AAT. The model refuse the classification request and responded with a vague and ambiguous reply.
Appendix D Extended Experiment Results
In this section, we illustrate more experiment results. Figure 8 provides a visual illustration of the average bias score distributions across various social groups and stereotype dimensions for 8 LLMs, where blue denotes the Competence, green denotes Sociability, and red denotes Morality.
The results show that most models exhibit bias scores close to +1 for groups such as Disability, Overweight, and Mental illness. Since positive bias indicates that favorable descriptors are disproportionately assigned to advantaged groups, these findings suggest that disadvantaged groups are systematically deprived of positive semantic associations. Although the magnitude of bias varies across model families, the overall trend is consistent: GPT-4-turbo and the LLaMa series display relatively stable distributions close to neutrality but still skewed toward advantaged groups; GPT-4o shows larger fluctuations; the DeepSeek series produces more dispersed results and sometimes extreme values; Claude-3.7-sonnet and Gemini-2.5-pro fall in between, with Claude exhibiting particularly strong bias against Mental illness. Overall, these patterns indicate that large language models tend to reinforce existing social hierarchies across sensitive dimensions such as competence, sociability, and morality, allocating more positive descriptors to advantaged groups.
Figure 9 presents the distribution of emotional framings across different social groups for each LLM, where green denotes the proportion of Comedy, red denotes Tragedy, and blue denotes Neutrality.
The y-axis represents the percentage of each emotional category, with the total summing to 100% for each group. The results show that most models produce a certain proportion of neutral responses, whereas the DeepSeek series (R1 and V3) generate almost no neutral outputs, with results heavily concentrated in either "comedy" or "tragedy". This suggests that DeepSeek relies less on avoidance mechanisms in its alignment strategy and instead tends to provide more definitive judgments, even if such polarized expressions may reinforce group-related stereotypes. These results indicate implicit biases in LLM behaviors.

### v3

@@ Introduction
| (Low Morality) | Unprincipled, Unethical, Dishonest, Deceptive, Unaccountable, Corrupt, Unreliable, Duplicitous, Hypocritical, Negligent, Unscrupulous, Fraudulent, Deceitful, Manipulative, Unjust, Lawless, Self-serving, Exploitative, Opportunistic |
| Category | Words |
| Furniture | Table, Chair, Shelf |
| Vessel | Bottle, Plate, Cup, Box, Bag, Container |
| Tool | Pen, Key, Map, Coin, Wire, Pipe, Tool |
| Structure | Bridge, Window, Door, Frame, Fence |
| Nature | Road, Cloud, Stone, Hill, Path |
| Object | Book, Sheet, Lamp, Clock |
Data Template Design. For each task, we design 3 sets of data templates to minimize potential unintended biases that may be introduced by specific prompt formulations.
Data Generation. Finally, we perform automated construction by randomly sampling word combinations from the lexicon resources and inserting them into data templates to generate complete dataset. Specifically, for the WABT, we construct 10 paired combinations of and (e.g., African vs. American, Asian vs. American, etc.). For each combination, we randomly sample one pair of and group identifiers, and subsequently sample 5 and 5 attribute words from the lexicons corresponding to the three stereotype content dimensions. This sampling procedure is repeated 50 times for each combination. The sampled items are then combined with 3 data templates, resulting in a total of 4,500 instances. For the AAT, we randomly sample 500 combinations of group identifiers and neutral nouns , and combine them with 3 data templates, resulting in a total of 1,500 instances.
Experiment Results
Evaluated Models
Evaluation results on WABT
We input each instance into LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further compute the frequency counts of 4 specific combinations: , , , and . Based on these counts, we calculate the bias score, following our predefined computational formulas.
After computing the bias scores for all data, we calculate the average bias score for each model along each stereotype dimension. We then conduct one-sample t-tests to assess whether the mean bias scores significantly deviated from 0. The results include the number of valid responses (), mean bias score (Mean), standard deviation of the bias scores (Std), t-statistic (), and significance level () for each model across the three dimensions, as summarized in the table. In general, larger -values indicate stronger bias tendencies, while smaller -values provide greater statistical confidence in the existence of such biases. The detailed results are presented in Table 4.
| Models | Dim. | Mean | Std | |||
| Claude-3.7-sonnet | Com. | 1470 | -0.006 | 0.933 | -0.263 | 0.792 |
| Soc. | 1482 | 0.415 | 0.830 | 19.273 | .001 | |
| Mor. | 1465 | -0.012 | 0.966 | -0.493 | 0.622 | |
| DeepSeek-R1 | Com. | 178 | -0.246 | 0.891 | -3.672 | .001 |
| Soc. | 183 | 0.320 | 0.787 | 5.492 | .001 | |
| Mor. | 129 | 0.085 | 0.915 | 1.056 | 0.293 | |
| DeepSeek-V3 | Com. | 42 | 0.054 | 0.985 | 0.354 | 0.725 |
| Soc. | 17 | 0.180 | 0.945 | 0.763 | 0.456 | |
| Mor. | 23 | 0.130 | 0.991 | 0.617 | 0.544 | |
| Gemini-2.5-pro | Com. | 1351 | -0.057 | 0.923 | -2.252 | 0.025 |
| Soc. | 1398 | 0.367 | 0.792 | 17.329 | .001 | |
| Mor. | 1352 | -0.012 | 0.931 | -0.489 | 0.625 | |
| GPT-4o | Com. | 382 | -0.067 | 0.923 | -1.423 | 0.156 |
| Soc. | 400 | 0.275 | 0.885 | 6.203 | .001 | |
| Mor. | 419 | 0.176 | 0.956 | 3.770 | .001 | |
| GPT-4-turbo | Com. | 1256 | -0.063 | 0.953 | -2.345 | 0.019 |
| Soc. | 1205 | 0.341 | 0.835 | 14.150 | .001 | |
| Mor. | 1160 | 0.072 | 0.976 | 2.512 | 0.012 | |
| LlaMa-2-70B-Chat | Com. | 99 | -0.039 | 0.698 | -0.558 | 0.578 |
| Soc. | 113 | 0.256 | 0.657 | 4.124 | .001 | |
| Mor. | 78 | 0.128 | 0.798 | 1.408 | 0.163 | |
| LlaMa-3-70B-Instruct | Com. | 1156 | 0.098 | 0.901 | 3.681 | .001 |
| Soc. | 1130 | 0.246 | 0.854 | 9.682 | .001 | |
| Mor. | 1095 | 0.065 | 0.920 | 2.337 | 0.020 |
Evaluation Results on AAT
We input each data into 8 LLMs and obtain their respective responses. For each instance, we record the number of valid responses returned by the models. For each valid response, we further analyze the emotional framing chosen by the model—specifically, whether the response aligns more closely with a comedic or tragic interpretation.
Notably, a substantial portion of responses appear ambiguous or equivocal, indicating that the model does not make a clear choice between comedy and tragedy. We categorize such responses as Neutrality. We then compute the proportion of responses labeled as comedy, tragedy, and neutrality separately for cases where the social entity belongs to either or . The results are presented in Table 5. The results reveal substantial variation in emotional framing across different social groups. Certain groups, such as Disability, Overweight, and Mental illness, are consistently associated with higher proportions of tragedy across multiple models, indicating a potential bias toward negatively valenced portrayals. In contrast, groups such as Asian, Youth, and American are more frequently linked with comedy or neutrality, suggesting relatively less stereotypical or emotionally charged representations. Moreover, some models demonstrate particularly polarized patterns. For instance, DeepSeek-R1 and DeepSeek-V3 show overwhelmingly tragic framings across almost all groups, while LLaMa-2-70B-Chat produces a predominance of neutral responses, especially for marginalized identities.
| Models | ||||||
| C. | T. | N. | C. | T. | N. | |
| Claude-3.7-sonnet | 0.777 | 0.048 | 0.175 | 0.780 | 0.047 | 0.173 |
| DeepSeek-R1 | 0.265 | 0.735 | 0.000 | 0.212 | 0.788 | 0.000 |
| DeepSeek-V3 | 0.267 | 0.725 | 0.008 | 0.279 | 0.708 | 0.013 |
| Gemini-2.5-pro | 0.340 | 0.628 | 0.032 | 0.352 | 0.624 | 0.024 |
| GPT-4o | 0.267 | 0.725 | 0.008 | 0.279 | 0.708 | 0.013 |
| GPT-4-turbo | 0.532 | 0.465 | 0.003 | 0.452 | 0.547 | 0.001 |
| LLaMa-2-70B-Chat | 0.368 | 0.025 | 0.607 | 0.372 | 0.027 | 0.601 |
| LLaMa-3-70B-Instruct | 0.567 | 0.388 | 0.045 | 0.574 | 0.388 | 0.038 |
In-Depth Analysis of MIST
Finding-1: A consistently observed pervasive positive bias in sociability. One of the most prominent and consistent findings in the WABT is the widespread presence of positive bias in the Sociability dimension. With the exception of Claude-3.7-sonnet, whose bias score mean is close to 0, the majority of models exhibit a statistically significant positive Sociability bias (mean 0, p .001). Notably, Gemini-2.5-pro demonstrates the highest average bias score in Sociability (0.367) among all models, accompanied by a relatively low standard deviation (0.792), indicating a consistent tendency to attribute higher Sociability traits to a wide range of social groups. Such a tendency may stem from inherent biases in the training data, such as a preference for positive interpersonal interactions or idealized personality traits. Alternatively, it may reflect an inductive prior embedded in the model’s design, aimed at generating responses perceived as helpful, cooperative, or socially appropriate.
Finding-2: Multidimensional complexity of divergent bias patterns across dimensions. Bias in LLMs varies substantially across stereotype dimensions in terms of direction, magnitude, and statistical significance, and may even exhibit opposing patterns. For example, DeepSeek-R1 shows a significant positive bias in Sociability (0.320, p .001), while exhibiting a significant negative bias in Competence (–0.246, p .001). In contrast, its bias in Morality is not statistically significant (0.085, p 0.293). This contrast indicates that bias patterns are not synchronized across dimensions. Similar heterogeneity is observed in other models. GPT-4o demonstrates significant positive bias in both Sociability and Morality, while its bias in Competence is statistically non-significant, suggesting a relatively neutral stance in that dimension. Overall, these results show that bias in LLMs is not a monolithic phenomenon, but rather a multi-faceted, dimension-specific structure.
Finding-3: Emergent neutral response patterns in affective attribution. In the AAT, models are instructed to make a binary affective attribution between Comedy and Tragedy. However, several models spontaneously generate a non-trivial proportion of Neutrality responses that are not specified in the predefined output space, indicating attributional avoidance or uncertainty under certain social contexts. As shown in Table 5, substantial variation exists across models in the prevalence of neutral outputs. LLaMa-2-70B-Chat exhibits the highest concentration of neutral attributions, reaching 60.7% for groups and 60.1% for groups. In contrast, Claude-3.7-sonnet and LLaMa-3-70B-Instruct maintain comparatively low neutral rates (approximately 4%–17%), though they still display stable neutral tendencies within specific subgroups. This emergent neutrality may reflect two underlying mechanisms: (1) exposure to sensitive social group identifiers induces internal stereotype conflicts, leading to indecisive or avoidance-oriented attribution behavior; and (2) safety-oriented alignment objectives encourage models to proactively avoid emotionally sensitive judgments, favoring neutralized responses as a form of implicit safety regulation.
Finding-4: Implicit bias exhibits asymmetric and divergent patterns. Across all evaluated LLMs, FAR and UAR do not increase simultaneously, indicating that affective attribution bias does not follow an idealized dual-peak distribution. Instead, most models exhibit elevated bias on only one of the two metrics. For instance, Claude-3.7-sonnet and LLaMa-3-70B-Instruct achieve higher FAR scores of 77.7% and 56.7%, respectively, suggesting a stronger tendency to assign favorable affective attributions to advantaged groups. In contrast, DeepSeek-R1 reaches 78.8% on the UAR metric, reflecting a stronger tendency to assign unfavorable affective attributions to disadvantaged groups. These results indicate that although all models display group-level attribution bias at a global level, the mechanisms through which such biases are expressed differ across models. Some models are primarily characterized by positive amplification toward advantaged groups, whereas others exhibit negative amplification toward disadvantaged groups.
Related Work
ToM in LLMs. Recent advancements in LLM reasoning [44, 43, 17, 16] have spurred debate on their potential for an emergent ToM. Evaluating this capacity is now a pivotal goal for human-centered AI [19, 22, 26, 41, 7, 25]. However, scholarly assessments diverge. Some critics argue that high performance stems from flawed benchmarks or superficial statistical patterns rather than genuine reasoning [42, 33], while other research shows ToM can be robustly trained to generalize [27]. A critical failure arises when LLMs misapply learned societal stereotypes to individuals.
Stereotype Content Model. SCM explains stereotypes along Warmth and Competence [11], with Warmth refined into Sociability and Morality [21], highlighting ambivalent prejudice [9, 13, 6]. Recent work shows that LLMs reproduce SCM-consistent patterns: despite generally positive tone, group descriptions align with SCM dimensions [20, 34], defaulting to white, healthy, middle-aged male representations while exhibiting semantic shifts for other groups [4, 37]. This work [30] analyses further indicate that Warmth and Competence remain dominant evaluative axes and that bias extends beyond binary labels, with implications for education and hiring [2, 45].
Conclusion
In this paper, we design a framework MIST to evaluate implicit social biases in LLMs by framing them as failures of ToM. We identify issues in existing evaluation methods, which often rely on direct-query tasks susceptible to social desirability effects and fail to capture the multi-dimensional nature of stereotypes. To address these, our method includes two main strategies: (1) reconceptualizing bias through the multi-dimensional Stereotype Content Model, and (2) developing the Word Association Bias Test and the Affective Attribution Test as indirect tasks to elicit latent biases. These strategies enable our framework to probe for biases along distinct psychological dimensions and bypass the models’ explicit bias-avoidance mechanisms, improving the detection of subtle, structural stereotype patterns.